# Relative monodromy of ramified sections on abelian schemes

### v3

@@ 1 Introduction
Let $S$ be a regular, irreducible, quasi-projective variety defined over $\mathbb{C}$. An abelian scheme $\phi\colon\mathcal{A}\to S$ over $\mathbb{C}$ of relative dimension $g$ can be seen as family of $g$-dimensional complex abelian varieties $\{\mathcal{A}_{s}\}_{s\in S(\mathbb{C})}$ (the fibres of $\phi$) that “varies algebraically accordingly to $\phi$” over the parameterizing space $S(\mathbb{C})$. Each fibre $\mathcal{A}_{s}$ can be analytically identified with a complex torus $\mathbb{C}^{g}/\Lambda_{s}$. Locally, on simply connected subsets $U\subset S(\mathbb{C})$ it is possible to define a holomorphic period map $\mathfrak{P}:U\to\operatorname{Lie}(\mathcal{A})^{2g}$ where $\operatorname{Lie}(\mathcal{A})$ is the complex Lie algebra bundle of $\mathcal{A}(\mathbb{C})$. The period map associates to any point $s\in U$ a $\mathbb{R}$-basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$ and in general it cannot be globally defined on $S(\mathbb{C})$. The obstruction on the existence of the analytic continuation of $\mathfrak{P}$ along a loop around $s\in S(\mathbb{C})$ is measured by the monodromy group $\textrm{Mon}(\mathcal{A},s)$. Given an algebraic section of the abelian scheme, which is a $\mathbb{C}$-morphism $\sigma\colon S\to\mathcal{A}$, again locally on simply connected open subsets $U^{\prime}\subset S(\mathbb{C})$ it is possible to lift $\sigma$ through the exponential map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$ to a holomorphic map $\log_{\sigma}\colon U^{\prime}\to\operatorname{Lie}(\mathcal{A})$. This map is called an abelian logarithm of the section $\sigma$. On a common open set $V\subset S(\mathbb{C})$ where both $\log_{\sigma}$ and $\mathfrak{P}$ are defined, we can express $\log_{\sigma}(s)\in\mathbb{C}^{g}$ in terms of the basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$. The coordinates form a real-analytic function $\beta_{\sigma}\colon V\to\mathbb{R}^{2g}$ which is nowadays called the Betti map (associated to $\sigma$) and that turned out to be of crucial importance in the theory of unlikely intersections. Such map was implicitly considered by Manin in [25], to get a proof of Mordell’s conjecture in the function field case. The aforementioned abelian logarithm $\log_{\sigma}$ is in general defined only locally, thus it is interesting to measure the obstruction to the existence of a global abelian logarithm: the simultaneous monodromy of the period map and abelian logarithm of $\sigma$ gives rise to the monodromy group of the section $M_{\sigma}$. The knowledge of $\textrm{Mon}(\mathcal{A},s)$ and the monodromy group of the logarithm of a section gives relevant information on the structure of the abelian scheme and on the section, respectively: in fact, no non-zero periods can be globally defined when the abelian scheme has no fixed part (see [16, Lemma 5.6]) and the monodromy of a logarithm is trivial only when the section is the zero-section of the scheme. It is usually important to understand more on the structure of monodromy groups. In [9] an [27] the authors construct examples where the monodromy group $\textrm{Mon}(\mathcal{A},s)$ is of infinite index in the arithmetic group attached to its Zariski closure, i.e. thin. Also in [30] it is pointed out that thin groups often arise in the context of monodromy groups. Given a finitely generated group in $\textrm{GL}_{n}(\mathbb{Z})$, usually it’s not too difficult to compute its Zariski closure, but deciding if the group is thin can be very subtle and difficult. Thus, in general it is very interesting to determine under which hypotheses the monodromy groups are big in their Zariski closures.
@@
We investigate the obstruction that prevents the existence of a globally defined abelian logarithm along loops which induce a trivial monodromy action on the period map. This obstruction is measured by the relative monodromy group $M^{\textrm{rel}}_{\sigma}$. It is easy to check that $M^{\textrm{rel}}_{\sigma}$ is a $\mathbb{Z}$-submodule of $\mathbb{Z}^{2g}$, so it is interesting to understand how its rank is related to $\sigma$. The group $M_{\sigma}^{\textrm{rel}}$ gives relevant information on the section, in particular its non-triviality should be related to the property of $\sigma$ of being non-torsion, and its rank should give information on the minimal group subscheme containing the image of $\sigma$.
In the case of non-isotrivial elliptic schemes Corvaja and Zannier proved in [10] that when $\sigma$ is a non-torsion section the group $M^{\textrm{rel}}_{\sigma}$ is non-trivial and moreover it is isomorphic to $\mathbb{Z}^{2}$ (so it has full rank). In the same paper, they provided an example showing that a priori, if we just limit ourselves to group theoretic arguments, the relative monodromy group could be smaller than expected (even trivial) when compared to the Zariski closures of non-relative monodromy groups. Hence, in this setting there are some groups which cannot arise as monodromy groups of sections. Moreover, we point out that the second author of the present paper in [33] studied the relative monodromy for products of two elliptic schemes. In [32] he also gives an explicit proof of the results of Corvaja and Zannier.
In this paper we generalize the work of [10] on the relative monodromy $M^{\textrm{rel}}_{\sigma}$, focusing on ramified sections instead of non-torsion sections. The notion of ramified section is formally introduced below but it essentially means that $\sigma:S\to\mathcal{A}$ doesn’t come from a base-change of an abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ whose modular map is unramified.
Let $\mathfrak{A}_{g,\ell}\to\mathbb{A}_{g,\ell}$ be a universal family with some fixed $-\ell$-structure and without locally constant parts; in order to simplify the notation, in the rest of the paper we drop the subscript $\ell$ and simply denote the family by $\mathfrak{A}_{g}\to\mathbb{A}_{g}$. Moreover, in order to make the following definition meaningful, we also need to assume that the base $S$ is maximal in the sense that on any compactification $\overline{A}\to\overline{S}$ the fibers over $\overline{S}\setminus S$ are all singular.
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme with modular map $p:S\to\mathbb{A}_{g}$. We say that a section $\sigma:S\to\mathcal{A}$ is unramified (over $\mathbb{A}_{g}$) if:
  • $(i)$
    There exist an unramified morphism $q^{\prime}:S^{\prime}\to p(S)$ and an abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ such that $\mathcal{A}=\mathcal{A}^{\prime}\times_{S^{\prime}}S$ and the following diagram is commutative:
[…]
  • $(ii)$
    There exists a section $\sigma^{\prime}:S^{\prime}\to\mathcal{A}^{\prime}$ with the following property: let $q^{*}\sigma^{\prime}\colon S\to\mathcal{A}$ be the natural section induced by $\sigma^{\prime}\circ q$, then $\sigma=q^{*}\sigma^{\prime}$.
A section $\sigma\colon S\to\mathcal{A}$ is said to be ramified (over $\mathbb{A}_{g}$) if it is not unramified.
Under the extra hypotheses that the abelian scheme is without fixed part (i.e. it has trivial Chow trace) and that the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite [^1] In the literature the is no uniform consensus on the precise notion of generically finite morphism. For us a morphism of schemes $f\colon X\to Y$ is generically finite if there exists a dense open $V\subseteq Y$ such that the restriction $f\colon f^{-1}(V)\to V$ is a finite morphism. (onto its image) we prove the result stated below. Notice that such condition on the modular map is often referred as maximal variation in moduli.
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is non-trivial.
Our techniques seem unsuitable to prove the same statement for unramified non-torsion sections. However, when it is possible to apply [26, Main Theorem] on the finiteness of the group of holomorphic sections for Kuga families, we extend 1.2 for all non-torsion sections. In particular, this happens when the modular map is generically finite onto a weakly-special subvariety of $\mathbb{A}_{g}$.
In addition, we provide a precise result on the size of $M^{\textrm{rel}}_{\sigma}$ in some special cases:
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section and the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
In a similar way as above 1.3 can be extended for all non-torsion sections under some additional hypothesis. Note that the action of $\textrm{Mon}(\mathcal{A})$ is irreducible for instance when $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$.
1.3 partially settles the following general conjecture:
@@
After the first version of this paper (which contained slightly weaker results) was made public, we privately received the preprint [2] of Y. André who partially addresses 1.4, including some of our cases. Under certain hypotheses, including the requirement that the monodromony of periods is not thin, he shows that $M^{\textrm{rel}}_{\sigma}$ is of maximal rank. See Section 3.5 for more details and an explicit example in which the monodromy of periods is thin and $M^{\textrm{rel}}_{\sigma}$ is of maximal rank.
@@
Our theorem has immediate implications on the arithmetic properties of the sections of abelian schemes, in particular on some properties of the Betti map. On one hand periods and abelian logarithms in general cannot be globally defined, but on the other hand it is also true that they can be always be globally defined after going to the universal cover $\widetilde{S}$ of the parameter base $S$, so the Betti map is also globally defined on $\widetilde{S}$. Let us denote by $S_{*}\to S$ the minimal unramified cover where periods are globally defined and by $S_{\sigma}\to S_{*}$ the minimal unramified cover where abelian logarithm of $\sigma$ is globally defined: then, $S_{\sigma}\to S_{*}$ is also the minimal unramified cover of $S_{*}$ where the Betti map is globally defined and we have the following situation
$\widetilde{S}\to S_{\sigma}\to S_{*}\to S\,.$
The map $S_{*}\to S$ is a Galois cover whose Galois group is $\textrm{Mon}(\mathcal{A},s)$, while the map $S_{\sigma}\to S$ is a Galois cover with Galois group $M_{\sigma}$. Our main theorem determines properties of the Galois group $M_{\sigma}^{\textrm{rel}}$ of the Galois cover $S_{\sigma}\to S_{*}$, giving precise information about the global definition of the Betti map which is useful to know for several applications.
@@
The Betti map is used in the work [29] of Pila and Zannier together with a result of Pila and Wilkie in transcendental Diophantine geometry (see [28]) to give a new proof of the Manin-Mumford conjecture. Such a strategy is referred as “the Pila-Zannier method”; in the case of abelian schemes over $\overline{\mathbb{Q}}$ it works in the following way: the goal is to have a control on the distribution of torsion values of $\sigma$ (i.e. elements of $\sigma^{-1}(\mathcal{A}_{\operatorname{tor}})$) with bounded heights. First one gives lower bounds on the number of such torsion values using Galois conjugates and the height inequality of Dimitrov-Gao-Habegger. On the other hand the Betti map $\beta_{\sigma}$ transforms the torsion values in rational points of a definable set in the $o$-minimal structure $\mathbb{R}_{\operatorname{an},\exp}$. At this point a “Pila-Wilkie type” result gives a control on the rational points on the transcendental part of such definable set. Still we have no information on the rational points in the algebraic part of the definable set, but here the crucial point is to use some functional transcendence theorem of “Ax-Schanuel type” to control the algebraic part of the definable set. Let us take a closer look at the last step involving the functional transcendence arguments: a classical way to tackle this step is by using a result of André (see [1, Theorem 3]). It says that the coordinates of the abelian logarithm $\log_{\sigma}$ are pairwise algebraically independent over the extension of $\mathbb{C}(S)$ generated by the coordinates of the period map $\mathfrak{P}$. Such result has often very strong implications, for instance in [14], thanks to such result, it is shown that the algebraic part of the definable set constructed with the Pila-Zannier method is empty.
@@
The approach of our proofs was originally inspired by [10], and it is based on the interpretation of the obstruction to globally define a logarithm as a Galois cohomology class with respect to the monodromy action of the fundamental group of $S(\mathbb{C})$ on $\mathbb{Z}^{2g}$. The new key ideas consist of using the Lefschetz hyperplane theorem for quasi-projective varieties (see [19, 2.2]) to restrict our reasoning to generic curves on the base. At this point we exploit the properties of the trace operator with respect to the pullback of the section in the case of finite morphisms, but as opposed to [10] we don’t need to construct abelian schemes with finite Mordell-Weil group. [^2] In a previous version of the paper we considered only Kuga families so that we could apply [26, Main Theorem].
For the proof of 1.2 we proceed in the following way: we construct an intermediate abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ which factorizes the universal diagram induced by the modular map $p:S\to\mathbb{A}_{g}$. This construction is carried out by means of topological techniques in such a way that $\mathcal{A}^{\prime}\to S^{\prime}$ is a finite unramified base change of the universal family. Then we find a suitable curve inside $S(\mathbb{C})$ which preserves the complete monodromy action of the fundamental group of $S(\mathbb{C})$ on periods and logarithms. The curve is obtained after iterating the aforementioned Lefschetz theorem and turns out to be a Stein space. This last property allows us to construct holomorphic sections of the unramified family which come from a Galois cohomology class describing the obstruction of the global logarithms. The proof then follows by a contradiction argument: we assume the triviality of $M_{\sigma}^{\textrm{rel}}$, then by using the previous apparatus and the properties of trace and pull-back operators, we deduce that $\sigma$ should be unramified. This proves the non-triviality of $M_{\sigma}^{\textrm{rel}}$.
@@
Overview of the paper.
In Section 2 we define all the main objects and we collect the results coming from the general theory. In particular, we give a detailed account of the various monodromy groups. In Section 3.2 we carry out the proof of 1.2 after some technical results about “the invariance” of our main theorem (Section 3.1). Section 3.3 is devoted to the proof of the corollary. In Section 3.4 and Section 3.5 we provide several examples of abelian schemes with sections for which we are able to estimate the relative monodromy. Finally, Section 4 is devoted to some useful applications of our results.
Acknowledgments and remarks.

### v4

@@ 1 Introduction
Notation and Background.
We establish the following notation, which will remain fixed throughout this paper. Let $S$ be a regular, irreducible, quasi-projective variety over $\mathbb{C}$, and let $\phi\colon\mathcal{A}\to S$ be an abelian scheme over $\mathbb{C}$ of relative dimension $g$. By definition, this means that $\mathcal{A}$ is a proper, smooth group $\mathbb{C}$-scheme over $S$ equipped with a zero section $\sigma_{0}\colon S\to\mathcal{A}$, and whose fibers $\mathcal{A}_{s}$ are abelian varieties of dimension $g$ for every $s\in S$. The generic fiber of $\mathcal{A}$ is denoted by $A$ and is an abelian variety over the complex function field $\mathbb{C}(S)$. Since any rational section of $\phi\colon\mathcal{A}\to S$ can be extended to a global section (see for instance [20, Proposition 6.2]), there is a group isomorphism between the rational points $A(\mathbb{C}(S))$ and $\Sigma(S)$ which denotes the set of (algebraic) sections of $\phi$. We recall that the fixed part of $A$ (sometimes called the “constant part”) is the image of the $\overline{\mathbb{C}(S)}/\mathbb{C}$-trace of $A\otimes\overline{\mathbb{C}(S)}$ (see for instance [9]). The fixed part is essentially the maximal abelian subcheme of $A$ defined over the base field $\mathbb{C}$, hence “constant” when spread over $S$. We say that the abelian scheme $\mathcal{A}$ is without fixed part if the fixed part of $A$ is $\{0\}$.
We denote with $\mathbb{A}_{g,\ell}$ the moduli space of $g$-dimensional principally polarized complex abelian varieties endowed with a $-\ell$-structure, with $\ell\geq 3$. Since $\mathbb{A}_{g,\ell}$ is a fine moduli space, it comes with a universal family $\mathfrak{A}_{g,\ell}\to\mathbb{A}_{g,\ell}$. To simplify notation, we shall omit the subscript $\ell$ in what follows, writing instead $\mathfrak{A}_{g}\to\mathbb{A}_{g}$. Furthermore, up to replacing $\mathcal{A}\to S$ with a base change by a finite étale morphism $S^{\prime}\to S$, by [16, Proof of Theorem B.1 Devissages $(iv)-(vi)$], we can assume that the abelian scheme carries principal polarization and it has $-\ell$-structure. Therefore, we fix a modular morphism $p\colon S\to\mathbb{A}_{g}$ induced by $\phi$, and denote its image by $T=p(S)$.
@@
For any complex variety $X$ we denote by $X(\mathbb{C})$ the associated analytification in the sense of [26, Exposé XII]. If $X$ is regular, then $X(\mathbb{C})$ is a complex manifold with the associated locally euclidean topology. We denote with $\Sigma^{\operatorname{an}}$ the sheaf over $S(\mathbb{C})$ of holomorphic sections of $\phi$ whereas $\Sigma(S)\subset\Sigma^{\operatorname{an}}(S(\mathbb{C}))$ is the abelian group of (global) algebraic $\mathbb{C}$-sections. Each fiber $\mathcal{A}_{s}(\mathbb{C})$ can be analytically identified with a complex torus $\mathbb{C}^{g}/\Lambda_{s}$. Locally, on simply connected subsets $U\subset S(\mathbb{C})$ it is possible to define a holomorphic period map $\mathfrak{P}:U\to\operatorname{Lie}(\mathcal{A})^{2g}$ where $\operatorname{Lie}(\mathcal{A})$ is the complex Lie algebra bundle of $\mathcal{A}(\mathbb{C})$. The period map associates to any point $s\in U$ a basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$ and in general it cannot be globally defined on $S(\mathbb{C})$. The obstruction on the existence of the analytic continuation of $\mathfrak{P}$ is measured by the monodromy group (of $\mathcal{A}$), $\textrm{Mon}(\mathcal{A})\subseteq\text{Sp}_{2g}(\mathbb{Z})$. The group $\textrm{Mon}(\mathcal{A})$ is often addressed also as monodromy of periods. Given an algebraic section $\sigma\colon S\to\mathcal{A}$ of the abelian scheme, locally on simply connected open subsets $U^{\prime}\subset S(\mathbb{C})$ it is possible to lift $\sigma$ through the exponential map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$ to a holomorphic map $\log_{\sigma}\colon U^{\prime}\to\operatorname{Lie}(\mathcal{A})$. This map is called an abelian logarithm of the section $\sigma$. On a common open set $V\subset S(\mathbb{C})$ where both $\log_{\sigma}$ and $\mathfrak{P}$ are defined, we can express $\log_{\sigma}(s)\in\mathbb{C}^{g}$ in terms of the basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$. The coordinates form a real-analytic function $\beta_{\sigma}\colon V\to\mathbb{R}^{2g}$ which is nowadays called the Betti map (associated to $\sigma$) and that turned out to be of crucial importance in the theory of unlikely intersections. Such map was implicitly considered by Manin in [32], to get a proof of Mordell’s conjecture in the function field case. The abelian logarithm $\log_{\sigma}$ is in general defined only locally, thus it is interesting to measure the obstruction to the existence of a global abelian logarithm: the simultaneous monodromy of the period map and $\log_{\sigma}$ gives rise to the monodromy group of the section, $M_{\sigma}\subseteq\textrm{SL}_{2g+1}(\mathbb{Z})$.
The non-triviality of $\textrm{Mon}(\mathcal{A})$ and $M_{\sigma}$ gives relevant information on the structure of the abelian scheme and on the section, with respect to the fixed part. In fact: no non-zero periods can be globally defined (so in particular $\textrm{Mon}(\mathcal{A})\neq\{0\}$) when the abelian scheme has no fixed part (see [21, Lemma 5.6]). In addition, if $M_{\sigma}=\begin{pmatrix}\mathrm{Mon}(\mathcal{A})&0\\ 0&1\end{pmatrix}$ then $\sigma$ is contained in the fixed part of $\mathcal{A}$ (see 3.10).
A deeper understanding of monodromy group structures, going beyond simply establishing their non-triviality, is often essential. For example, $\textrm{Mon}(\mathcal{A})\subseteq\textrm{Sp}_{2g}(\mathbb{Z})$ measures the “modular complexity” of $\phi\colon\mathcal{A}\to S$, and a key question is whether it is thin, i.e. of infinite index in $G(\mathbb{Z})$, where $G$ denotes its Zariski closure. Works such as [12] and [34] exhibit important cases where $\textrm{Mon}(\mathcal{A})$ is thin. While thin monodromy groups appear to be more common in certain settings, non-thin examples also arise (see, e.g., [37] or [18, Section 2.1] for a discussion). For a finitely generated subgroup of $\textrm{GL}_{n}(\mathbb{Z})$, determining its Zariski closure is usually computationally feasible. However, verifying whether the group is thin can be a delicate and highly non-trivial problem.
@@
We investigate the obstruction that prevents the existence of a globally defined abelian logarithm along loops which induce a trivial monodromy action on the period map. This obstruction is measured by the relative monodromy group (of the section) $M^{\textrm{rel}}_{\sigma}\subseteq\mathbb{Z}^{2g}$. The group $M_{\sigma}^{\textrm{rel}}$ gives relevant information on the section, in particular its non-triviality is supposed be related to the property of $\sigma$ of being non-torsion, and its rank should give information on the minimal group subscheme containing the image of $\sigma$.
In the case of non-isotrivial elliptic schemes over a curve Corvaja and Zannier proved in [13] that when $\sigma$ is a non-torsion section the group $M^{\textrm{rel}}_{\sigma}$ is non-trivial and moreover it is isomorphic to $\mathbb{Z}^{2}$ (so it has full rank). In the same paper, they provided an example showing that a priori, if we just limit ourselves to group theoretic arguments, the relative monodromy group could be smaller than expected (even trivial) when compared to the relative group arising from Zariski closures of non-relative monodromy groups. Hence, in this setting there are some groups which cannot arise as monodromy groups of sections. We point out that the second author of the present paper in [40] gives an explicit proof of the results of Corvaja and Zannier.
In this paper we generalize the work of [13] on the relative monodromy $M^{\textrm{rel}}_{\sigma}$, focusing on ramified sections instead of non-torsion sections. The concept of a ramified section is formally defined below. Intuitively, a section $\sigma\colon S\to\mathcal{A}$ is ramified if it does not arise from a base change of an abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ whose modular map is unramified and factors $p$. The main technical difficulty arises when $T$ is non-normal, requiring a more careful definition.
Let $\nu\colon T^{\nu}\to T$ denote the normalization of $T$, and let $\mathfrak{A}_{g,T^{\nu}}\to T^{\nu}$ be the base change of the universal family to $T^{\nu}$. By the universal property of normalization, the modular map $p\colon S\to T$ factors uniquely through $\nu$, yielding a dominant map $p^{\nu}\colon S\to T^{\nu}$. Let $L$ be the maximal unramified extension [^1] Let $X$ be an integral, normal scheme. Recall that an extension $L$ of $K(X)$ is said unramified if the normalization $X^{\prime}\to X$ of $X$ in $L$ is an unramified morphism. of $\mathbb{C}(T^{\nu})$ in $\mathbb{C}(S)$ and let $S^{\operatorname{ur}}$ be the normalization of $T^{\nu}$ in $L$. By the universal property of the normalization (in a field extension) we get a factorization
[…]
where $q^{\prime}$ is unramified, $q$ is dominant and $S^{\operatorname{ur}}$ is normal. By taking the fiber product over $T^{\nu}$ we fix once for all the following commutative diagram
[…]
@@
With the above notation fixed we say that an algebraic section $\sigma:S\to\mathcal{A}$ is unramified (over $S$) if there exists an algebraic section $\sigma^{\prime}:S^{\operatorname{ur}}\to\mathcal{A}^{\prime}$ such that $\sigma=q^{*}\sigma^{\prime}$, where
$q^{*}\sigma^{\prime}\colon S(\mathbb{C})$$\to$ $\mathcal{A}(\mathbb{C})$
$s$ $\mapsto$ $(\mathfrak{q}^{-1}_{s}\circ\sigma^{\prime}\circ q)(s)$
and $\mathfrak{q}_{s}:\mathcal{A}_{s}\xrightarrow{\cong}\mathcal{A}_{q(s)}$ is the isomorphism induced by the fiber product, for any $s\in S(\mathbb{C})$. An algebraic section $\sigma\colon S\to\mathcal{A}$ is said to be ramified (over $S$) if it is not unramified. We say that $\sigma$ is unramified [resp. ramified] over a Zariski open subset $U\subseteq S$ if $\sigma_{|U}$ is an unramified [resp. ramified] section of the restricted abelian scheme $\mathcal{A}_{|U}\to U$.
In LABEL:{cor:divisible_ram} we show that every torsion section is unramified. This fact is noted in [13] for elliptic schemes over a curve, though we were unable to find a proof or reference for higher-dimensional cases. Of particular interest are settings where unramified sections are necessarily torsion, as this condition is equivalent to say that every abelian scheme obtained via an unramified base change of $\mathfrak{A}_{g,T}\to T$ has zero Mordell-Weil rank. For example, Shioda [38] proves that unramified sections of elliptic schemes over a curve are always torsion, while Mok and To [33] establish the vanishing of the Mordell-Weil rank for unramified finite extension Kuga families (without locally constant part). However, in general torsion sections form a proper subset of unramified sections, as in 2.4 we construct a non-torsion unramified section.
In the following, we consider abelian schemes without fixed part. Additionally, we assume that the modular map $p:S\to T\subseteq\mathbb{A}_{g}$ is generically finite [^2] For us a morphism of schemes $f\colon X\to Y$ is generically finite if there exists a dense open $V\subseteq Y$ such that the restriction $f\colon f^{-1}(V)\to V$ is a finite morphism. (onto its image), this assumption is often referred as maximal variation in moduli. When $\mathcal{A}\to S$ is simple and $S$ is a curve, maximal variation in moduli is equivalent to being without fixed part. We prove the result stated below.
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that the modular map $p:S\to T\subseteq\mathbb{A}_{g}$ is generically finite. Let $U_{p}\subseteq S$ be the Zariski open subset on which $p$ is finite and flat. Then for any section $\sigma\colon S\to\mathcal{A}$ ramified over $U_{p}$, the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is non-trivial.
When $\phi\colon\mathcal{A}\to S$ is an elliptic scheme over a curve, we may always assume that the modular map $p$ is finite and flat after performing a base change. In this case we have $U_{p}=S$, and the results of [13] are recovered. An interesting case where our result holds, is given by the family of Jacobians of hyperelliptic curves described in [4, Section 8.1]. Our techniques seem unsuitable to prove the same statement for all non-torsion sections or for a more general modular map $p$. However, in the particular case when $p$ is finite and flat and all unramified sections of $\mathcal{A}\to S$ are torsion, then the claim 1.2 clearly holds for all non-torsion sections. For instance, by [33, Main Theorem] such special situation occurs when $p$ is finite and $\mathfrak{A}_{g,T}\to T$ is a Kuga family without locally constant part. In particular, $T$ may be chosen to be a totally geodesic (i.e. weakly-special) subvariety of $\mathbb{A}_{g}$.
Regarding the relation between the rank of $M^{\textrm{rel}}_{\sigma}$ and the geometric properties of $\sigma$ we prove the following:
@@
Under the same hypotheses as in 1.2, if the (linear) action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then for any section $\sigma\colon S\to\mathcal{A}$ ramified over $U_{p}$, the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
Also 1.3 can be extended for all non-torsion sections under some additional hypothesis. For instance the action of $\textrm{Mon}(\mathcal{A})$ is irreducible when $T$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$ (see [4, 10.2.6 Lemma]) or when the abelian scheme is simple. Hence our 1.3 holds for all sections, which are ramified outside a Zariski closed subset, of a simple abelian scheme without fixed part over a curve $S$. Our results suggest the following conjecture:
@@
1.3 proves 1.4 when the modular map is finite and surjective onto $\mathbb{A}_{g}$. Notice that such a hypothesis prevents our abelian scheme from having any non-trivial proper group-subscheme. For the same reason, in this case the scheme automatically is without fixed part. Moreover, the second author of the present paper in [41] proved 1.4 in the case of products of two elliptic schemes over a curve.
After the initial version of this paper (containing somewhat weaker results) was circulated, Y. André kindly sent us a preprint [3], which addresses certain special cases of 1.4 when $g^{\prime}=g$, overlapping with parts of our work. His approach is fundamentally Hodge-theoretic, and his main hypothesis on the abelian scheme is the non-thinness of $\mathrm{Mon}(\mathcal{A})$, in contrast to our condition on the modular map $p$. A more detailed comparison is provided in Section 4.4.
@@
Our theorem has immediate implications on the arithmetic properties of the sections of abelian schemes, in particular on some properties of the Betti map. On one hand periods and abelian logarithms in general cannot be globally defined, but on the other hand it is also true that they can be always be globally defined after going to the universal cover $\widetilde{S}$ of the base $S(\mathbb{C})$, so the Betti map is also globally defined on $\widetilde{S}$. Let us denote by $S_{*}\to S(\mathbb{C})$ the minimal topological cover where periods are globally defined and by $S_{\sigma}\to S_{*}$ the minimal topological cover where abelian logarithm of $\sigma$ is globally defined: then, $S_{\sigma}\to S_{*}$ is also the minimal topological cover of $S_{*}$ where the Betti map is globally defined and we have the following situation
$\widetilde{S}\to S_{\sigma}\to S_{*}\to S(\mathbb{C})\,.$
The map $S_{*}\to S(\mathbb{C})$ is a Galois cover whose Galois group is $\textrm{Mon}(\mathcal{A},s)$, while the map $S_{\sigma}\to S(\mathbb{C})$ is a Galois cover with Galois group $M_{\sigma}$. Our main theorem determines properties of the Galois group $M_{\sigma}^{\textrm{rel}}$ of the Galois cover $S_{\sigma}\to S_{*}$, giving precise information about the global definition of the Betti map which is useful to know for several applications.
@@
The Betti map is used in the work [36] of Pila and Zannier together with a result of Pila and Wilkie in transcendental Diophantine geometry (see [35]) to give a new proof of the Manin-Mumford conjecture. Such a strategy is referred as “the Pila-Zannier method”; in the case of abelian schemes over $\overline{\mathbb{Q}}$ it works in the following way: the goal is to have a control on the distribution of torsion values of $\sigma$ (i.e. elements of $\sigma^{-1}(\mathcal{A}_{\operatorname{tor}})$) with bounded heights. First one gives lower bounds on the number of such torsion values using Galois conjugates and the height inequality of Dimitrov-Gao-Habegger. On the other hand the Betti map $\beta_{\sigma}$ transforms the torsion values in rational points of a definable set in the $o$-minimal structure $\mathbb{R}_{\operatorname{an},\exp}$. At this point a “Pila-Wilkie type” result gives a control on the rational points on the transcendental part of such definable set. Still we have no information on the rational points in the algebraic part of the definable set, but here the crucial point is to use some functional transcendence theorem of “Ax-Schanuel type” to control the algebraic part of the definable set. Let us take a closer look at the last step involving the functional transcendence arguments: a classical way to tackle this step is by using a result of André (see [2, Theorem 3]). It says that the coordinates of the abelian logarithm $\log_{\sigma}$ are pairwise algebraically independent over the extension of $\mathbb{C}(S)$ generated by the coordinates of the period map $\mathfrak{P}$. Such result has often very strong implications, for instance in [17], thanks to such result, it is shown that the algebraic part of the definable set constructed with the Pila-Zannier method is empty.
@@
Our proofs develop the ideas introduced in [13], interpreting the obstruction to defining a global logarithm as a Galois cohomology class relative to the monodromy action of $\pi_{1}(S(\mathbb{C}))$ on $\mathbb{Z}^{2g}$. The key innovation involves applying the Lefschetz hyperplane theorem for quasi-projective varieties (see [24, Part II, 1.2, Theorem]) to reduce the problem to generic curves in the base variety. We then utilize properties of the trace operator with respect to the pullback of sections (in the case of finite flat morphisms). Unlike the approach in [13], our method does not require the construction of abelian schemes with finite Mordell-Weil groups.
For the proof of 1.2 we proceed in the following way: we construct an intermediate abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ which factorizes the universal diagram induced by the modular map $p:S\to T$. This construction is carried out by means of topological techniques in such a way that $\mathcal{A}^{\prime}\to S^{\prime}$ is a finite unramified base change (after normalization) of the universal family. Then we find a suitable curve inside $S(\mathbb{C})$ which preserves the complete monodromy action of the fundamental group of $S(\mathbb{C})$ on periods and logarithms. The curve is obtained after iterating the aforementioned Lefschetz theorem and turns out to be a Stein space. This last property allows us to construct holomorphic sections of the unramified family which come from a Galois cohomology class describing the obstruction of the global logarithms. The proof then follows: we assume the triviality of $M_{\sigma}^{\textrm{rel}}$, then by using the previous apparatus and the properties of trace and pull-back operators, we deduce that $\sigma$ should be unramified. This proves the non-triviality of $M_{\sigma}^{\textrm{rel}}$ for ramified sections (over $U_{p}$).
@@
Acknowledgmentss.